OpenStudy (anonymous):

What is the inverse of the function f(x) = x +3? a. h(x) =1/3x + 3 b. h(x) = 1/3x – 3 c. h(x) = x – 3 d. h(x) = x + 3

OpenStudy (anonymous):

@madhu.mukherjee.946

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

Begin with the function: f(x) = x +3 I think of it as y = x +3 Interchange the y and x. x = y + 3 Solve for y. y = x - 3 y = h(x) h(x) = x - 3

OpenStudy (madhu.mukherjee.946):

thats what i told

Directrix (directrix):

>thats what i told The answer you posted but not the work.